it takes a runner 18 seconds to run a distance of 264 feet. What is the runners speed in mph?
Answer by jim_thompson5910(35256)

18 seconds * (1 min/60 seconds)*(1 hr/60 min) = 0.005 hours

264 feet * (1 mile/5280 feet) = 0.05 miles

-------------------------------------------------------

d = rt

0.05 = r*0.005

0.05/0.005 = r

10 = r

r = 10

So the runner's speed is 10 miles per hour.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------